# Semantic-Aware Resource Management for C-V2X Platooning via Multi-Agent Reinforcement Learning

### v1

@@ Abstract
This paper presents a semantic-aware multi-modal resource allocation (SAMRA) for multi-task using multi-agent reinforcement learning (MARL), termed SAMRAMARL, utilizing in platoon systems where cellular vehicle-to-everything (C-V2X) communication is employed. The proposed approach leverages the semantic information to optimize the allocation of communication resources. By integrating a distributed multi-agent reinforcement learning (MARL) algorithm, SAMRAMARL enables autonomous decision-making for each vehicle, channel assignment optimization, power allocation, and semantic symbol length based on the contextual importance of the transmitted information. This semantic-awareness ensures that both vehicle-to-vehicle (V2V) and vehicle-to-infrastructure (V2I) communications prioritize data that is critical for maintaining safe and efficient platoon operations. The framework also introduces a tailored quality of experience (QoE) metric for semantic communication, aiming to maximize QoE in V2V links while improving the success rate of semantic information transmission (SRS). Extensive simulations has demonstrated that SAMRAMARL outperforms existing methods, achieving significant gains in QoE and communication efficiency in C-V2X platooning scenarios.
@@ I-A Background
With the rapid development of intelligent transportation systems (ITS) [1, 2, 3], ensuring safe and efficient transportation is increasingly crucial [4, 5, 6]. A key component of ITS is platooning systems, where multiple autonomous vehicles travel closely together to enhance traffic flow and safety [7, 8, 9, 10, 11]. In these systems, a designated platoon leader (PL) manages formation, while platoon members (PMs) maintain coordinated speed and distance. Effective intra-platoon and inter-platoon communication (platoon-to-platoon or platoon-to-infrastructure) is essential for optimizing overall efficiency and safety.
Integrating cellular vehicle-to-everything (C-V2X) communication is imperative for achieving effective communication. C-V2X supports two main types: vehicle-to-vehicle (V2V), which allows vehicles to share cooperative awareness messages (CAMs) for synchronized movement, and vehicle-to-infrastructure (V2I), enabling communication with base stations for traffic and safety information [12, 13, 14, 15, 16, 17, 18]. These modalities facilitate timely responses to dynamic traffic conditions, enhancing the safety and efficiency of platooning systems [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
However, increased connectivity and communication demands in C-V2X platooning introduce complexities in managing network resources, which are critical for reliable communication in autonomous driving [30, 31, 32, 33]. The frequency of information exchange among vehicles directly impacts their ability to react to obstacles, underscoring the need for efficient resource management [34, 35, 36, 37, 38].
A promising approach to these challenges is semantic communication [39], which focuses on transmitting meaningful information rather than raw data. This enhances data transmission efficiency in C-V2X systems by ensuring only relevant information is shared [40], improving decision-making and reducing unnecessary data transfer.
As task complexity rises, there is a shift from unimodal to multi-modal tasks that integrate various data types [41, 42, 43, 44]. This integration fosters a richer understanding of the environment, enhancing network robustness and flexibility [45].
Traditional centralized resource management schemes often face inefficiencies and high signaling overhead due to their reliance on global information, particularly under dynamic channel conditions [46, 47].
To address these challenges, we propose a distributed semantic-aware multi-modal resource allocation (SAMRA) framework leveraging multi-agent reinforcement learning (MARL), termed SAMRAMARL [^1] The source code has been released at: https://github.com/qiongwu86/Semantic-Aware-Resource-Management-for-C-V2X-Platooning-via-Multi-Agent-Reinforcement-Learning. By employing MARL, vehicles can make decentralized decisions based on local observations, reducing reliance on centralized control while enhancing scalability and adaptability to dynamic network conditions [48].
@@ I-B Related Work and Motivation
Resource allocation in platooning systems has garnered significant attention within intelligent transportation systems (ITS). Platooning involves multiple autonomous vehicles traveling in coordination, necessitating efficient communication between the platoon leader (PL), other platoon members (PMs), and infrastructure like base stations [49]. Various studies have explored traditional methods to address these challenges.
For instance, Guo et al. [50] proposed a joint optimization approach for LTE-V2V radio resource allocation and vehicle control parameters to enhance platoon stability. Hong et al. [51] designed a framework using relays and adaptive distributed model predictive control (DMPC) to improve safety in failure scenarios. Wang et al. [52] developed a two-step resource allocation strategy optimizing platoon formation and power control through branch and bound methods. Wen et al. [53] focused on optimizing inter-vehicle communication topology in LTE-V2V networks. However, the dynamic nature of channel conditions complicates effective resource management due to uncertainties in estimating channel state information (CSI).
@@
In response, semantic communication has gained traction. Unlike traditional methods that focus on raw data transmission, semantic communication prioritizes the meaning behind the conveyed information. This approach can enhance coordination and safety in platooning systems [59]. For instance, Bourtsoulatze et al. [60] developed a method for efficient image transmission under low signal-to-noise ratio (SNR) conditions, while Huang et al. [61] used generative adversarial networks for semantic image compression.
@@ I-C Contributions
In this work, we present a distributed semantic-aware multi-modal resource allocation (SAMRA) algorithm based on MARL, termed SAMRAMARL, tailored for platooning systems. The primary contributions of our research are summarized as follows:
@@
    Simulation results demonstrate that our SAMRAMARL algorithm significantly outperforms existing algorithms in terms of QoE and SRS.
The rest of the paper is structured as follows: Section II introduces the system model and the formulated problem of maximizing QoE and SRS; Sections III present the proposed SAMRA algorithm; Section IV provides and discusses simulation results; and Section V concludes the paper.
@@ II System Model
As shown in Fig. 1, we consider a C-V2X communication network that integrates a BS with $N$ platoons of automated vehicles. Each platoon $n$ consists of ${M_{n}}$ vehicles, where $n=\left\{{1,2,\ldots,N}\right\}$, and the first vehicle in each platoon is designated as the PL, which coordinates intra-platoon communication and handles inter-platoon communication with BS. The set of vehicles in each platoon is denoted as $m=\left\{{1,\ldots,M}\right\}$. The communication system supports two types of communication: intra-platoon (V2V) and inter-platoon (V2I). In intra-platoon communication, vehicles within a platoon share information directly via V2V links to synchronize movement and optimize traffic efficiency. Inter-platoon communication occurs between the PL of each platoon and the BS over V2I links, where control and safety messages are exchanged to maintain coordination across platoons. The communication system utilizes orthogonal frequency-division multiplexing (OFDM), with bandwidth divided into orthogonal subchannels, each with a bandwidth of $W$. Both V2V and V2I communications take place over these subchannels, allowing efficient bandwidth management and interference mitigation.
This system supports two communication modes: single-modal tasks and multi-modal tasks. For single-modal communication, the deep semantic communication (DeepSC) model is employed, which is designed for textual data exchange [65]. For multi-modal tasks, the multi-user deep semantic communication (MU-DeepSC) model is utilized, supporting both text and image data transmission [66]. In inter-platoon communication, vehicles primarily engage in single-modal communication, exchanging textual semantic information with the BS. The textual data is converted into semantic symbols using the DeepSC model. For instance, if a vehicle detects a sudden braking event, it could transmit the corresponding semantic information as a warning message to the BS, which then disseminates this information to the relevant platoons. In intra-platoon communication, the use of multi-modal tasks is determined by the number of PMs. If the number of PMs is even, vehicles engage in multi-modal communication, where each pair of vehicles employs the MU-DeepSC model to exchange both textual and image-based semantic information. If the number of PMs is odd, multi-modal communication is applied to pairs of vehicles, while the remaining vehicle communicates use single-modal transmission. In contrast, the inter-platoon (V2I) communication mode depends on the number of PLs communicating with BS. When the number of PLs is even, all PLs engage in multi-modal communication, transmitting both text and image data. When the number of PLs is odd, most of the PLs communicate employ multi-modal communication, while one remaining PL uses single-modal communication, focusing solely on textual data. For example, within a platoon, one vehicle captures an unfamiliar traffic sign while another vehicle inputs a text query to inquire about the sign’s meaning. Both pieces of semantic information are sent to the PL, who, after processing, provides an accurate interpretation of the sign, like “This sign indicates a pedestrian crossing ahead.”
@@
II-A Semantic Transceivers
II-A1 DeepSC Transmitter
@@
where $s{e_{{\alpha_{T}}}}(\cdot)$ represents the semantic encoding function utilizing a Bi-directional long short-term memory (Bi-LSTM) network, parameterized by ${\alpha_{T}}$. This encoder extracts meaningful semantic features from the input text. Subsequently, $c{h_{{\beta_{T}}}}\left(\cdot\right)$, the channel encoder, compresses and maps the semantic features into transmitted symbols suitable for communication over the physical channel. The combined operation ensures that only the essential, compressed semantic information is transmitted, optimizing resource usage in the V2X environment.
II-A2 MU-DeepSC Transmitter
@@
where $s{e_{{\alpha_{I}}}}(\cdot)$ is the image semantic encoder (based on a deep convolutional neural network, such as ResNet-101), and $c{h_{{\beta_{I}}}}(\cdot)$ is the channel encoder responsible for compressing the extracted semantic features into symbols. The text input follows a similar encoding process as described for the DeepSC model. Both the image-transmitted symbols ${X_{I}}$ and the text-transmitted symbols ${X_{T}}$ are transmitted together or sequentially, depending on the network’s transmission protocol. This approach ensures that both image and textual semantic information are conveyed efficiently, maintaining the integrity of multi-modal data while optimizing bandwidth and reducing redundancy.
II-A3 Transmission Model
After semantic and channel encoding, the transmitted symbols ${X_{T}}$ (textual) and ${X_{I}}$ (image) are sent over the wireless channels in the platooning system. We consider two types of communication: intra-platoon communication (V2V) and inter-platoon communication (V2I), both subject to channel effects and interference. To present the wireless transmission model with a clearer distinction between V2V (intra-platoon) and V2I (inter-platoon) communication, we define the binary decision variable ${\rho_{n,k}}\in\{0,1\}$. This variable indicates the choice of communication type for platoon $n$ over subchannel $k$. Specifically, when ${\rho_{n,k}}=1$, PL uses subchannel $k$ for V2V communication to broadcast messages within its own platoon; otherwise, the PL uses subchannel $k$ for V2I communication to communicate with BS. After semantic and channel encoding, the transmitted symbols ${X_{T,n}}$ (text) and ${X_{I,n}}$ (image) from the PL are transmitted over the wireless channel in the platooning system. The received signals, interference, and SINR are modeled separately for V2V and V2I communication, taking into account the choice of ${\rho_{n,k}}$.
In V2V communication, vehicles exchange information through direct communication links within the same platoon. Let ${X_{T,n}}$ and ${X_{I,n}}$ represent the transmitted symbols for text and image data from the PL of platoon $n$. The received signals at vehicle $m$ within platoon $n$ over subchannel $k$ are expressed as follows: for text transmission:
$Y_{T,n,m}[k]=\rho_{n,k}h_{n,m}[k]X_{T,n}[k]+I_{T,n,m}[k]+{\chi_{T,n,m}},$ (3)
@@ II-A3 Transmission Model
$Y_{I,n,m}[k]=\rho_{n,k}h_{n,m}[k]X_{I,n}[k]+I_{I,n,m}[k]+{\chi_{I,n,m}},$ (4)
where $\rho_{n,k}h_{n,m}[k]$ is the channel gain between the PL of platoon $n$ and vehicle $m$ over subchannel $k$, $I_{T,n,m}[k]$ and $I_{I,n,m}[k]$ represent the interference from other platoons transmitting on the same subchannel, and ${\chi_{T,n,m}}$ and ${\chi_{I,n,m}}$ denote the additive white Gaussian noise (AWGN) for text and image transmissions, respectively. The corresponding signal-to-interference-plus-noise ratio (SINR) for text and image transmissions are given by: for text transmission:
$\text{SINR}_{T,n,m}[k]=\frac{\rho_{n,k}p_{T,n}[k]h_{n,m}[k]}{I_{T,n,m}[k]+\sigma^{2}},$ (5)
@@
$\text{SINR}_{I,n,m}[k]=\frac{\rho_{n,k}p_{I,n}[k]h_{n,m}[k]}{I_{I,n,m}[k]+\sigma^{2}},$ (6)
@@
where $(1-\rho_{n,k})h_{n,BS}[k]$ is the channel gain between the PL of platoon $n$ and the BS over subchannel $k$ when $\rho_{n,k}=0$. $I_{T,\text{BS}}[k]$ denotes the interference from other platoons using the same subchannel for V2I communication, and $\chi_{T,\text{BS}}$ represents the noise at the BS. The SINR for inter-platoon communication for text transmission is:
$\text{SINR}_{T,\text{BS}}[k]=\frac{(1-\rho_{n,k})p_{T,n}[k]h_{n,\text{BS}}[k]}{I_{T,\text{BS}}[k]+\sigma^{2}}.$ (8)
The total interference for intra-platoon and inter-platoon communications is modeled as follows: For text transmission:
${I_{T,n,m}}[k]=\sum\limits_{n^{\prime}\neq n}{{\rho_{n^{\prime},k}}{\beta_{n^{\prime},k}}}{p_{T,n^{\prime}}}[k]{h_{n^{\prime},m}}[k],$ (9)
for image transmission:
${I_{I,n,m}}[k]=\sum\limits_{n^{\prime}\neq n}{{\rho_{n^{\prime},k}}{\beta_{n^{\prime},k}}}{p_{I,n^{\prime}}}[k]{h_{n^{\prime},m}}[k],$ (10)
where ${{\beta_{n^{\prime},k}}}$ indicates whether subchannel $k$ is allocated to platoon ${n^{\prime}}$. where $\rho_{n^{\prime},k}$ indicates that subchannel $k$ is used for V2V communication or V2I communication by other platoons $n^{\prime}$. and ${{p}_{T,n^{\prime}}}[k]$ and ${{p}_{I,n^{\prime}}}[k]$ are the transmit powers for text and image data, respectively.
II-A4 Receiver Model

### v2

@@ Abstract
Semantic communication transmits the extracted features of information rather than raw data, significantly reducing redundancy, which is crucial for addressing spectrum and energy challenges in 6G networks. In this paper, we introduce semantic communication into a cellular vehicle-to-everything (C-V2X)-based autonomous vehicle platoon system for the first time, aiming to achieve efficient management of communication resources in a dynamic environment. Firstly, we construct a mathematical model for semantic communication in platoon systems, in which the DeepSC model and MU-DeepSC model are used to semantically encode and decode unimodal and multi-modal data, respectively. Then, we propose the quality of experience (QoE) metric based on semantic similarity and semantic rate. Meanwhile, we consider the success rate of semantic information transmission (SRS) metric to ensure the fairness of channel resource allocation. Next, the optimization problem is posed with the aim of maximizing the QoE in vehicle-to-vehicle (V2V) links while improving SRS. To solve this mixed integer nonlinear programming problem (MINLP) and adapt to time-varying channel conditions, the paper proposes a distributed semantic-aware multi-modal resource allocation (SAMRA) algorithm based on multi-agent reinforcement learning (MARL), referred to as SAMRAMARL. The algorithm can dynamically allocate channels and power and determine semantic symbol length based on the contextual importance of the transmitted information, ensuring efficient resource utilization. Finally, extensive simulations have demonstrated that SAMRAMARL outperforms existing methods, achieving significant gains in QoE, SRS, and communication delay in C-V2X platooning scenarios.
@@ I-A Background
With the rapid development of intelligent transportation systems (ITS) [1], ensuring safe and efficient transportation is increasingly crucial. A key component of ITS is platooning systems, where multiple autonomous vehicles travel closely together to enhance traffic flow and safety [2]. In these systems, a designated platoon leader (PL) manages formation, while platoon members (PMs) maintain coordinated speed and distance. Effective intra-platoon and inter-platoon communication (platoon-to-platoon or platoon-to-infrastructure) is essential for optimizing overall efficiency and safety.
Integrating cellular vehicle-to-everything (C-V2X) communication is imperative for achieving effective communication. C-V2X supports two main types: vehicle-to-vehicle (V2V), which allows vehicles to share cooperative awareness messages (CAMs) for synchronized movement, and vehicle-to-infrastructure (V2I), enabling communication with base stations (BSs) for traffic and safety information [3]. These modalities facilitate timely responses to dynamic traffic conditions, enhancing the safety and efficiency of platooning systems. However, increased connectivity and communication demands in C-V2X platooning introduce complexities in managing network resources, which are critical for reliable communication in autonomous driving [4]. The frequency of information exchange among vehicles directly impacts their ability to react to obstacles, underscoring the need for efficient resource management [5, 6].
A promising approach to these challenges is semantic communication [7], which focuses on transmitting extracted features (meaningful information) rather than raw data. This enhances data transmission efficiency in C-V2X systems by ensuring only relevant information is shared [8], improving decision-making and reducing unnecessary data transfer.
As task complexity rises, there is a shift from unimodal to multi-modal tasks that integrate various data types [9]. This integration fosters a richer understanding of the environment, enhancing network robustness and flexibility [10]. Moreover, traditional centralized resource management schemes often face inefficiencies and high signaling overhead due to their reliance on global information, particularly under dynamic channel conditions [11, 12, 13].
To address these challenges, we propose a distributed semantic-aware multi-modal resource allocation (SAMRA) framework leveraging multi-agent reinforcement learning (MARL), termed SAMRAMARL [^1] The source code has been released at: https://github.com/qiongwu86/Semantic-Aware-Resource-Management-for-C-V2X-Platooning-via-Multi-Agent-Reinforcement-Learning. By employing MARL, vehicles can make decentralized decisions based on local observations, reducing reliance on centralized control while enhancing scalability and adaptability to dynamic network conditions [14, 15].
@@ I-B Related Work and Motivation
Resource allocation in platooning systems has garnered significant attention within intelligent transportation systems (ITS). Platooning involves multiple autonomous vehicles traveling in coordination, necessitating efficient communication between the platoon leader (PL), other platoon members (PMs), and infrastructure like BSs [16]. Various studies have explored traditional methods to address these challenges.
For instance, Guo et al. [17] proposed a joint optimization approach for LTE-V2V radio resource allocation and vehicle control parameters to enhance platoon stability. Hong et al. [18] designed a framework using relays and adaptive distributed model predictive control (DMPC) to improve safety in failure scenarios. Wang et al. [19] developed a two-step resource allocation strategy optimizing platoon formation and power control through branch and bound methods. Wen et al. [20] focused on optimizing inter-vehicle communication topology in LTE-V2V networks. However, the dynamic nature of channel conditions complicates effective resource management due to bias in estimating channel state information (CSI).
@@
In response, unlike traditional methods that focus on raw data transmission, semantic communication prioritizes the meaningful features behind the conveyed information. This approach can enhance coordination and safety in platooning systems [26]. For instance, Bourtsoulatze et al. [27] developed a method for efficient image transmission under low signal-to-noise ratio (SNR) conditions, while Huang et al. [28] used generative adversarial networks for semantic image compression.
@@ I-C Contributions
In this work, we present a distributed SAMRA algorithm based on MARL, termed SAMRAMARL, tailored for platooning systems. The primary contributions of our research are summarized as follows:
@@
    Simulation results demonstrate that our SAMRAMARL algorithm significantly outperforms existing algorithms in terms of QoE, SRS, and communication delay.
The rest of the paper is structured as follows: Section II introduces the system model and the formulated problem of maximizing QoE and SRS; Sections III present the proposed SAMRA algorithm; Section IV provides and discusses simulation results; and the paper is concluded in Section V.
@@ II System Model
As shown in Fig. 1, we consider a C-V2X communication network that integrates a BS with $N$ platoons of autonomous vehicles. The set of platoons is denoted as $\mathcal{N}=\left\{{1,2,\ldots,N}\right\}$, $N\in\mathbb{N}$. Each platoon $n$ consists of ${M_{n}}$ vehicles, where $n\in\mathcal{N}$, and the first vehicle in each platoon is designated as the PL, which coordinates intra-platoon communication and handles inter-platoon communication with BS. The set of vehicles in platoon $n$ is denoted as $\mathcal{M}_{n}=\left\{{1,\ldots,M_{n}}\right\}$, $M_{n}\in\mathbb{N}$. The communication system supports two types of communication: intra-platoon (V2V) and inter-platoon (V2I). In intra-platoon communication, the PL broadcasts semantic messages to its followers, and PMs transmit semantic symbols to the PL via V2V links. This guarantees the platoon string stability and ensures that all the platoon members are aware of each other’s kinematics and decision-making. Inter-platoon communication occurs between the PL of each platoon and the BS over V2I links, where control and safety messages are exchanged to maintain coordination across platoons. Orthogonal frequency division multiplexing (OFDM) is utilized for communication. The total system bandwidth is divided into $K$ orthogonal subchannels with bandwidth $W$, and the set of subchannels is denoted as $\mathcal{K}=\left\{{1,\ldots,K}\right\}$, $K\in\mathbb{N}$. Both V2V and V2I communications take place over these subchannels, allowing efficient bandwidth management and interference mitigation.
@@
In the autonomous vehicle platoon, text and image data serve as the core modalities: textual data directly conveys explicit instructions or descriptions, while visual data enhances environmental awareness. By integrating these two modalities, the system can make appropriate decisions and maintain safety. Therefore, this system supports two communication modes: single-modal tasks and multi-modal tasks. For single-modal communication, the deep semantic communication (DeepSC)  [^2] DeepSC is not the only approach, but it is relatively mature and has been widely verified in terms of semantic compression and anti-noise performance. model is employed, which is designed for textual data exchange [32]. For multi-modal tasks, the multi-user deep semantic communication (MU-DeepSC) model is utilized, supporting both text and image data transmission [33]. Note that when both models generate semantic symbols, they will prioritize according to the contextual importance. For example, the text of the emergency braking warning will be encoded as high-priority semantic symbols. In inter-platoon communication, vehicles primarily engage in single-modal communication, exchanging textual semantic information with the BS. The textual data is converted into semantic symbols using the DeepSC model. For instance, if a vehicle detects a sudden braking event, it could transmit the corresponding semantic information as a warning message to the BS, which then disseminates this information to the relevant platoons. In intra-platoon communication, the use of multi-modal tasks is determined by the number of PMs. If the number of PMs is even, vehicles engage in multi-modal communication, where each pair of vehicles employs the MU-DeepSC model to exchange both textual and image-based semantic information. If the number of PMs is odd, multi-modal communication is applied to pairs of vehicles, while the remaining vehicle communicates use single-modal transmission. For example, within a platoon, one vehicle captures an unfamiliar traffic sign while another vehicle inputs a text query to inquire about the sign’s meaning. Both pieces of semantic information are sent to the PL, who, after processing, provides an accurate interpretation of the sign, like “This sign indicates a pedestrian crossing ahead”.
@@
II-A Semantic Communication Model
II-A1 Semantic Transceivers
@@
where $s{e_{{\alpha_{T}}}}(\cdot)$ represents the semantic encoding function utilizing a Bi-directional long short-term memory (Bi-LSTM) network, parameterized by ${\alpha_{T}}$. This encoder extracts meaningful semantic features from the input text. Subsequently, $c{h_{{\beta_{T}}}}\left(\cdot\right)$, the channel encoder, compresses and maps the semantic features into transmitted symbols suitable for communication over the physical channel. The combined operation ensures that only the essential, extracted features full of semantic information is transmitted, optimizing resource usage in the V2X environment.
@@
where $s{e_{{\alpha_{I}}}}(\cdot)$ is the image semantic encoder (based on a deep convolutional neural network, such as ResNet-101), and $c{h_{{\beta_{I}}}}(\cdot)$ is the channel encoder responsible for translating the extracted semantic features into symbols. The text input follows a similar encoding process as described for the DeepSC model. Both the image-transmitted symbols ${X_{I}}$ and the text-transmitted symbols ${X_{T}}$ are transmitted together or sequentially, depending on the network’s transmission protocol. This approach ensures that both image and textual semantic information are conveyed efficiently, maintaining the integrity of multi-modal data while optimizing bandwidth and reducing redundancy.
II-A2 Transmission Model
After semantic and channel encoding, the transmitted symbols ${X_{T}}$ (textual) and ${X_{I}}$ (image) are sent over the wireless channels in the platooning system. We consider a distributed resource allocation system where the platoons choose resources autonomously. However, hidden node problems or decision delays may lead to subchannel conflicts and hence interference. For example, two distant platoons choose the same sub-channel at the same time because they cannot sense each other. Alternatively, there may be a transient overlap in the sub-channel selection decisions of different platoons in dynamic environments.
The system consists of two types of communication: intra-platoon communication (V2V) and inter-platoon communication (V2I), both subject to channel effects and interference. To present the wireless transmission model with a clearer distinction between V2V (intra-platoon) and V2I (inter-platoon) communication, we define the binary decision variable ${\rho_{n,k}}\in\{0,1\}$. This variable indicates the choice of communication type for platoon $n$ over subchannel $k\in\mathcal{K}$. Specifically, when ${\rho_{n,k}}=1$, PL uses subchannel $k$ for V2V communication to broadcast messages within its own platoon; otherwise, the PL uses subchannel $k$ for V2I communication to communicate with BS. After semantic and channel encoding, the transmitted symbols ${X_{T,n}}$ (text) and ${X_{I,n}}$ (image) from the PL of platoon $n$ are transmitted over the wireless channel in the platooning system. The received signals, interference, and SINR are modeled separately for V2V and V2I communication, taking into account the choice of ${\rho_{n,k}}$.
In V2V communication, vehicles exchange information through direct communication links within the same platoon. The received signals at vehicle $m\in\mathcal{M}_{n}$ within platoon $n$ over subchannel $k$ are expressed as follows: for text transmission:
$Y_{T,n,m}[k]=\rho_{n,k}h_{n,m}[k]X_{T,n}[k]+I_{n,m}[k]+{\chi_{n,m}},$ (3)
@@ II-A2 Transmission Model
$Y_{I,n,m}[k]=\rho_{n,k}h_{n,m}[k]X_{I,n}[k]+I_{n,m}[k]+{\chi_{n,m}},$ (4)
where $\rho_{n,k}h_{n,m}[k]$ is the channel gain between the PL of platoon $n$ and vehicle $m$ over subchannel $k$, $I_{n,m}[k]$ represents the interference from other platoons transmitting on the same subchannel, and $\chi_{n,m}$ denotes the additive white Gaussian noise (AWGN), respectively. $I_{n,m}[k]$ is modeled as follows:
${I_{n,m}}[k]=\sum\limits_{n^{\prime}\neq n}{{\rho_{n^{\prime},k}}{\beta_{n^{\prime},k}}}{p_{n^{\prime}}}[k]{h_{n^{\prime},m}}[k],$ (5)
where ${{\beta_{n^{\prime},k}}}$ indicates whether subchannel $k$ is allocated to platoon ${n^{\prime}}$, $\rho_{n^{\prime},k}$ indicates that subchannel $k$ is used for V2V communication or V2I communication by other platoons $n^{\prime}$, and ${{p}_{n^{\prime}}}[k]$ is the transmit power, respectively. The corresponding signal-to-interference-plus-noise ratio (SINR) for text and image transmissions are given by: for text transmission:
$\text{SINR}_{T,n,m}[k]=\frac{\rho_{n,k}p_{T,n}[k]h_{n,m}[k]}{I_{n,m}[k]+\sigma^{2}},$ (6)
@@
$\text{SINR}_{I,n,m}[k]=\frac{\rho_{n,k}p_{I,n}[k]h_{n,m}[k]}{I_{n,m}[k]+\sigma^{2}},$ (7)
@@
where $(1-\rho_{n,k})h_{n,BS}[k]$ is the channel gain between the PL of platoon $n$ and the BS over subchannel $k$ when $\rho_{n,k}=0$. $I_{T,\text{BS}}[k]$ denotes the interference from other platoons using the same subchannel for V2I communication, and $\chi_{T,\text{BS}}$ represents the noise at the BS. $I_{T,\text{BS}}[k]$ is modeled as follows:
${I_{T,BS}}[k]=\sum\limits_{n^{\prime}\neq n}{{(1-\rho_{n^{\prime},k})}{\beta_{n^{\prime},k}}}{p_{T,n^{\prime}}}[k]{h_{n^{\prime},BS}}[k],$ (9)
The SINR for inter-platoon communication is:
$\text{SINR}_{T,n,\text{BS}}[k]=\frac{(1-\rho_{n,k})p_{T,n}[k]h_{n,\text{BS}}[k]}{I_{T,\text{BS}}[k]+\sigma^{2}}.$ (10)
II-A3 Receiver Model